A bottle holds 60L of a 55% saline solution. 20L are drained and replaced with another solution. Which reduces the solution of the mixture to 40% saline. What was the concentration of solution that was added?
HELP, PLEASE

Answer by lwsshak3(11628)

A bottle holds 60L of a 55% saline solution. 20L are drained and replaced with another solution. Which reduces the solution of the mixture to 40% saline. What was the concentration of solution that was added?
**
Amt of 55% solution remain after draining=60-20=40 L
amt of replacing solution=20 L
let x=% concentration of replacing solution
55%*40+x*20=40%*60
22+x*20=24
x*20=2
x=2/20=10
concentration of replacing solution=10%